OpenStudy (anonymous):

Would x^2+4/x+2 dx be the same thing as x^2+4(x+2)^-1dx?

OpenStudy (tkhunny):

Blech. No. Do you mean \(\dfrac{x^{2}+4}{x+2} = (x^{2}+4)(x+2)^{-1}\)? That's good, but you really have written nothing like that. Please remember your Order of Operations and try that again. Use parentheses where necessary. As far as the integral goes, simplify with long division and you'll have an easier time of it.

OpenStudy (anonymous):

long division?

OpenStudy (anonymous):

\[\int\limits \left( x ^{2}+4 \right)\left( -x-2 \right)dx\] like that?^

OpenStudy (tkhunny):

That is not long division. That is not what negative exponents mean. You need to do MUCH better than that. 

OpenStudy (triciaal):

your original yes it's the same